OpenStudy (anonymous):

I need help with my algebra homework. Please explain how to get the answers. Part A. Solve -np-90<30 for n Part B. Solve 5m-3h=45 for h

OpenStudy (tacotime):

A: -np-90<30 add 90 to both sides to get -np alone: -np<120 then divide by p: -n<120/p then multiply both sides by -1 (so you can get n not -n): n>120/p note that the less than sign changed direction. When ever you change the sign of both sides the inequality changes direction. This is just a rule. B:-5m-3h=45 move 5m with addition: -3h=45+3m then divide by -3: h = -(45+3m)/3

OpenStudy (anonymous):

Thank you